# Qizhe Zhang1,2  Aosong Cheng1  Ming Lu3  Zhiyong Zhuo1  Minqi Wang1 Jiajun Cao1  Shaobo Guo2  Qi She2  Shanghang Zhang1

### v1

@@ 3.3 Text-Visual Attention Dispersion
Figure 4: Concentration of [CLS] attention and last attention. The first row displays the input image and maps of [CLS] and the last attentions received by corresponding regions. The second row shows the CDF and the intensity distribution of [CLS] and the last attention. From multiple perspectives, it is evident that [CLS] attention is more concentrated than last attention.
In addition to the attention location distribution, we also analyze the distribution of attention intensity to further illustrate the lack of concentration in text-visual attention. The concentration here implies that a small number of tokens receive the majority of attention, while most tokens have low attention scores, indicating that the model identifies important tokens with high certainty. In row 1 of Fig. 4, we visualize the attention of the [CLS] token and the last output token on the visual tokens within the same image. It is apparent that [CLS] attention is highly focused, with only a few tokens receiving significant attention distributed in different positions like the aircraft nose, tail, runway, and sky. These tokens capture global information about the input image, and retaining them can accelerate VLM inference while preserving most of the performance. In contrast, the last output token attention (as well as other text tokens) is more dispersed, with multiple regions across the image receiving high attention, making it challenging to select important visual tokens during pruning. It is worth noting that in the last attention map, tokens with high attention in the [CLS] attention map receive little attention, while tokens with high attention scores are concentrated in the lower part of the image, including areas like the runway, grass, and even padding, which contain little valuable information. This aligns with the attention drift phenomenon discussed in the previous section.
In row 2 of Fig. 4, we first compute the Cumulative Distribution Function (CDF) of attention scores for visual tokens sorted by their attention scores, as shown in the left sub-figure. The CDF curve of [CLS] attention is significantly steeper than that of last attention, indicating that [CLS] attention is more concentrated. In the [CLS] attention distribution, the top 20% of visual tokens receive about 80% of total attention, compared to over 60% for last attention. In the remaining two sub-figures, we respectively show the average distribution of [CLS] and last attention intensity across $N$ samples. Most visual tokens receive minimal [CLS] attention, while last attention is more uniformly distributed, which means it contains more noise. The lack of attention dispersion in [CLS] attention suggests it is a more suitable indicator for guiding visual token pruning.
4 [CLS] Attention for Visual Token Pruning
Based on the above analysis of attention in VLMs, we propose FasterVLM, which uses [CLS] attention from the visual encoder as a more accurate indicator for visual token pruning. By removing redundant visual tokens before the LLM decoder, our approach could make VLM inference faster than methods that prune tokens within the LLM.
4.1 FasterVLM
Figure 5: Illustration of FasterVLM. We first re-rank image tokens using [CLS] attention from the visual encoder and prune the last R%. The remaining image tokens, after passing through the multi-modal projector, are combined with language instructions as input to the language model for response generation. Since redundant image tokens are removed before the language model, FasterVLM can make the inference of the entire VLM even faster than pruning within the language model.
To achieve more accurate visual token pruning, we propose FasterVLM, which leverages [CLS] attention from the visual encoder to rerank visual tokens and filter out the lower parts. Fig. 5 illustrates our FasterVLM design. At the output layer of the visual encoder (e.g. the penultimate layer of CLIP [45] in LLaVA-1.5 [32]), the attention from the [CLS] token ${\bm{z}}_{\text{[{CLS}]}}\in\mathbb{R}^{d}$ to other image tokens ${\bm{Z}}_{v}\in\mathbb{R}^{n\times d}$ is computed by:
${\bm{q}}_{\text{[{CLS}]}}={\bm{z}}_{\text{[{CLS}]}}{\bm{W}}_{Q},\quad{\bm{K}}_{v}={\bm{Z}}_{v}{\bm{W}}_{V},\quad{\bm{a}}_{\text{[{CLS}]}}=\text{Softmax}\left(\frac{{\bm{q}}_{\text{[{CLS}]}}{\bm{K}}_{v}^{T}}{\sqrt{d}}\right)$ (4)
where $n$ is the length of image tokens, $d$ is the size of hidden states, and ${\bm{W}}_{Q},{\bm{W}}_{V}$ is the transformation matrices for queries and keys at this layer. Using this [CLS] attention, we derive a dynamic threshold based on the computation budget $R$:
$\tau=\min\left\{\tau\mid\left|\{a\in{\bm{a}}_{\text{[{CLS}]}}\mid a\geq\tau\}\right|\leq n\times(1-R)\right\}$ (5)
In practice, $\tau$ is set as the $(1-R)$ percentile of $a_{\text{[{CLS}]}}$. With this threshold, we can prune unimportant tokens by:
${\bm{Z}}_{v}^{\prime}=\{{\bm{z}}_{v}^{i}\in{\bm{Z}}_{v}\mid{\bm{a}}_{\text{[{CLS}]}}^{i}\geq\tau\}$ (6)
The remaining image tokens, after passing through a multi-modal projector $g$, are concatenated with language instructions ${\bm{H}}_{q}$ and fed into the language model $f_{\phi}$ to generate the response ${\bm{X}}_{a}$:
${\bm{X}}_{a}=f_{\phi}([g({\bm{Z}}_{v}^{\prime});{\bm{H}}_{q}])$ (7)
Here, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes the concatenate operation. By performing token pruning before the LLM decoder, we dynamically adjust the number of visual tokens as input to the language model based on the actual computational budget, thus accelerating the VLM inference.
4.2 Theoretical Computation Reduction
Since the length of language instructions is uncertain and typically much shorter than the length of visual tokens, we only consider the FLOPs associated with visual tokens. In the language model of the VLM, assume $n$ is the number of visual tokens, $d$ is the hidden state size and $m$ is the intermediate size of FFN (with SwiGLU [49] activation in LLaMA [54]). For the prefill stage, the FLOPs of a single transformer layer can be estimated by $8nd^{2}+4n^{2}d+6ndm$, and the theoretical FLOPs reduction with a token reduction ratio $R$ (and $\hat{n}=(1-R)\cdot n$) is calculated as:
$F=1-\frac{8\hat{n}d^{2}+4\hat{n}^{2}d+6\hat{n}dm}{8nd^{2}+4n^{2}d+6ndm}=R+\frac{2n}{4d+2n+3m}(R-R^{2})$ (8)
Since the second term is relatively small, the overall FLOP reduction is close to $R$ but slightly higher. In the decode stage with KV cache, the attention complexity reduces to $\mathcal{O}(n)$, and the FLOPs for a transformer layer can be re-estimated by $8d^{2}+4nd+6ndm$. In this case, the theoretical FLOPs reduction is still close to $R$ but slightly lower. Moreover, existing hardware does not adapt well to the varying KV cache length during inference, so pruning before the LLM decoder (our FasterVLM implementation) actually results in more inference acceleration.
@@ 5 Experiments
In this section, we validate our method across various VLM architectures on comprehensive multi-modal benchmarks, including high-resolution image and video understanding tasks. We compare our approach with multiple existing methods and conduct an efficiency analysis.
@@ 5.1 Experimental Setup
Datasets. We conduct extensive experiments on 10 image-based multi-modal benchmarks, including common visual question answering tasks such as VQAv2 [15], GQA [20], VizWiz [16], ScienceQA-IMG [36], and TextVQA [50], as well as other multi-modal benchmarks such as POPE [29], MME [14], MMBench [35], MMBench-CN [35] and MM-Vet [62]. Additionally, we also experiment on 4 widely used video question answering benchmarks, including TGIF-QA [21], MSVD-QA [57], MSRVTT-QA [57], and ActivityNet-QA [63]. All experiments on these benchmarks follow the default settings and evaluation metrics. Details of each task are provided in the appendix.
Model architectures. We apply FasterVLM to various VLM architectures, including the classic LLaVA-1.5 [32] models with both 7B and 13B parameters, LLaVA-NeXT-7B [33] for high-resolution image inputs, and Video-LLaVA-7B [30] for video understanding. For all models, we follow the same inference settings as the original papers.
Comparison methods. We choose FastV [8], FitPrune [61], and SparseVLM [65] as comparison methods. All methods rely on text-visual attention in the LLM decoder for visual token pruning without additional training. FastV, as the initial approach, performs a one-time pruning after the second layer in the LLM. FitPrune derives an optimal pruning strategy based on predefined computational budgets, gradually pruning at each layer. SparseVLM, on the other hand, pre-selects text prompts used to guide pruning, aiming to reduce noise in text-visual attention.
@@ 5.2 Main results
Table 1: Performance comparison of different methods with LLaVA-1.5-7B across 10 benchmarks at reduction ratios ranging from 50% to 95%. For each reduction ratio, the best average performance is shown in bold. Our FasterVLM consistently achieves the best performance across all reduction ratios. Notably, even with 95% of visual tokens pruned, our method maintains 89.41% performance, demonstrating its effectiveness in accurately identifying the most critical visual information.
Method Reduction Ratio # Token VQAv2 GQA VizWiz SQA-IMG TextVQA POPE MME MMB MMB-CN MM-Vet Average
LLaVA-1.5-7B 0% 576 78.52 61.94 50.06 69.51 58.21 85.88 1506.47 64.69 58.08 31.30 100.00%
FastV 50% 288 77.67 60.05 50.53 68.96 58.25 82.45 1513.06 64.26 58.16 31.70 99.33%
FitPrune 78.41 61.70 50.04 69.16 58.26 85.37 1499.70 64.60 58.16 31.10 99.73%
SparseVLM 76.67 58.78 51.03 68.57 57.49 83.87 1458.79 63.14 56.87 31.50 98.52%
FasterVLM 77.86 60.64 50.45 68.37 57.90 86.20 1471.96 63.83 56.70 34.80 100.12%
FastV 75% 144 74.07 56.58 51.29 69.11 57.38 73.74 1463.39 64.00 57.22 28.60 95.80%
FitPrune 76.14 59.38 51.30 69.01 56.49 80.75 1472.86 63.92 57.65 28.40 97.22%
SparseVLM 72.76 55.11 51.46 69.36 55.99 77.57 1351.65 59.54 51.03 29.90 93.84%
FasterVLM 76.19 58.34 51.97 67.92 57.07 83.46 1433.76 62.54 57.13 34.20 98.75%
FastV 90% 58 65.38 51.20 51.84 69.81 54.75 57.30 1210.36 59.97 51.72 27.20 87.97%
FitPrune 62.76 49.96 50.85 68.22 50.35 53.81 1147.46 56.27 45.53 21.80 82.07%
SparseVLM 62.90 48.86 49.36 67.23 48.99 65.82 1030.61 49.05 35.40 18.60 78.13%
FasterVLM 71.92 54.91 53.01 68.91 55.28 75.85 1348.63 60.57 54.90 30.10 94.24%
FastV 95% 29 55.92 46.03 49.10 70.00 51.56 35.47 971.56 50.17 42.18 18.90 74.93%
FitPrune 52.39 43.60 48.61 68.32 46.75 31.17 855.21 39.69 29.98 18.00 67.64%
FasterVLM 66.75 51.51 52.67 69.56 53.09 67.24 1254.80 58.51 51.98 27.50 89.41%
We first apply FasterVLM to the widely used LLaVA-1.5 model and conduct a comprehensive comparison with existing methods. Tab. 1 shows the performance on the LLaVA-1.5-7B model as the visual token reduction ratio increases from 50% to 95%. We also plot the relationship between average performance and the visual token reduction ratio in Fig. 2. Note that SparseVLM includes a visual token recycling mechanism, which prevents reducing the visual token count below 95%. The results indicate that for all text-visual attention-based methods, model performance degrades noticeably as the reduction ratio increases, suggesting that text-based attention is not reliable in identifying the most important visual tokens. In contrast, our FasterVLM significantly mitigates this performance decline, retaining 89.41% of the original performance at a high reduction ratio of 95%, while substantially enhancing computational efficiency. This demonstrates the critical importance of [CLS] attention for effective visual token pruning.
We also validate our FasterVLM on the 13B version of LLaVA-1.5, with results shown in Tab. 3. Due to space constraints, we report only the average results across the 10 benchmarks. Detailed results can be found in the appendix. With a larger language model, the VLM achieves better performance on vision-language tasks, even with the same visual encoder. However, as the token pruning ratio increases, pruning methods based on text attention exhibit a similar, sharp decline in performance. In contrast, our method retains 88.35% of the original performance at a 95% reduction ratio, demonstrating that it achieves outstanding performance even with a larger language model.
5.3 FasterVLM with higher resolution
Table 2: Performance comparison of different methods with LLaVA-1.5-13B under different reduction ratios, with results reported as the average of 10 benchmarks. For each reduction ratio, the best average performance is shown in bold.
Method Reduction Ratio # Token Benchmark Average
LLaVA-1.5-13B 0% 576 100.00%
FastV 75% 144 97.16%
SparseVLM 94.32%
FasterVLM 97.43%
FastV 90% 58 90.26%
SparseVLM 85.02%
FasterVLM 93.68%
FastV 95% 29 80.53%
FasterVLM 88.35%
Table 3: Performance comparison of different methods with LLaVA-NeXT under different reduction ratios, with results reported as the average of 10 benchmarks. For each reduction ratio, the best average performance is shown in bold.
Method Reduction Ratio # Token Benchmark Average
LLaVA-NeXT-7B 0% 2880 100.00%
FastV 75% 720 97.37%
SparseVLM 94.95%
FasterVLM 98.73%
FastV 90% 290 89.24%
SparseVLM 88.65%
FasterVLM 93.55%
FastV 95% 145 77.43%
FasterVLM 88.85%

### v2

@@ 3.3 Text-visual attention dispersion
Figure 3: Analysis of text-visual attention dispersion. (a) Cumulative distribution of different attentions. In [CLS] attention, a few visual tokens absorb the majority of attention, while last attention is relatively dispersed. (b) Density distribution of different attentions. Unlike [CLS] attention, the density of last attention, both before and after the elimination of position embedding decay, is obviously high entropy and low peak. (c) Performance of pruning based on different attentions across various benchmarks. Last attention, with position embedding decay removed, achieves performance gains but still underperforms random pruning on some benchmarks due to the dispersion issue, while pruning based on [CLS] attention consistently takes the lead.
Given the negative impact of attention shift caused by rotary position embedding on pruning, an intuitive idea is to use text-visual attention devoid of position embedding decay as a basis for pruning. Unfortunately, although this practice eliminates positional bias, we find another phenomenon present both before and after the removal of position embedding, which we call attention dispersion. Specifically, we refer to last attention with identical position embedding for all visual tokens as equi last attention, and plot the cumulative distribution curves of different attentions in Fig. 3(a). It is observable that [CLS] attention is more concentrated, with the top 20% visual tokens absorbing about 80% attention. In contrast, text-visual attention, whether or not position embedding is removed, is spread across more visual tokens. We also illustrate the density distribution of different attentions in Fig. 3(b). Different from the extreme peak in [CLS] attention, the density of last attention both before and after removing position embedding shows high entropy and low peak, resembling a more uniform distribution. Here, highly concentrated attention indicates that the model identifies important tokens with high certainty, whereas a uniform distribution implies greater difficulty in selecting important tokens based on attention scores.
To verify this, we evaluate the performance of pruning based on different attentions across various benchmarks, as shown in Fig. 3(c). After eliminating the long-term decay brought by position embedding, pruning results improved across all benchmarks, further demonstrating the adverse impact of positional bias. However, it is worth noting that pruning based on last attention still lags behind random pruning on some benchmarks, especially the hallucination benchmark POPE. We attribute this gap partly to attention dispersion, which makes it difficult to identify important tokens with rich visual information based on attention. Pruning based on [CLS] attention consistently achieves optimal results across all benchmarks, motivating us to exploit visual cues for more effective token pruning in VLMs.
4 Exploiting Visual Cues for Token Pruning
Figure 4: Illustration of VisPruner. We begin by selecting a small portion of important tokens with rich information, based on the [CLS] attention from the visual encoder. For the remaining tokens, we progressively remove duplicates based on similarity, ultimately retaining another set of diverse tokens. These two parts complement each other, ensuring that the model maintains comparable performance even after a significant reduction of visual tokens, without relying on any additional training.
Based on the above analysis of attention in VLMs, we propose VisPruner, which exploits visual cues for more effective token pruning. We first select a small portion of important tokens with rich information based on the attention from the visual encoder in Sec. 4.1. Then, we retain another set of diverse tokens from the remaining ones based on their similarity as a complement in Sec. 4.2. These two parts together maintain the high performance of the model even after a significant reduction of visual tokens. The overall design of our VisPruner is illustrated in Fig. 4.
4.1 Attention-based important tokens
Visual inputs inherently possess redundancy, and the attention score is a widely used metric to assess token importance and information content [9, 60, 65, 55]. However, as analyzed before, attention within the language model, due to the presence of shift and dispersion phenomena, does not effectively identify truly important visual tokens. Therefore, in VisPruner, we utilize attention from the visual encoder to select important tokens. Specifically, for the attention matrix ${\bm{A}}$ calculated in Eq. 2, we first average across the attention heads. Then, for visual encoders like CLIP [43] that contain a [CLS] token, we take the first row of ${\bm{A}}$ as ${\bm{a}}_{v}\in\mathbb{R}^{n}$, representing the attention from the [CLS] token to other patch tokens, where $n$ is the length of visual tokens. For models without a [CLS] token (i.e. SigLIP [63]), we directly average the rows of ${\bm{A}}$ as the average attention each patch token receives from all other tokens. After getting the visual attention ${\bm{a}}_{v}$, we can derive a dynamic threshold $\tau$ based on the computation budget $R$:
$\tau=\min\left\{\tau\mid\left|\{a\in{\bm{a}}_{v}\mid a\geq\tau\}\right|\leq n\times(1-R\cdot r)\right\}$ (4)
Here, $r$ represents the ratio of important tokens within the total computation budget, which is a hyperparameter. In practice, $\tau$ is set as the $(1-R\cdot r)$ percentile of ${\bm{a}}_{v}$. With this threshold, we can select the important part ${\bm{Z}}_{v}^{imp}$ with rich information from all visual tokens:
${\bm{Z}}_{v}^{imp}=\{{\bm{z}}_{v}^{i}\in{\bm{Z}}_{v}\mid{\bm{a}}_{v}^{i}\geq\tau\}$ (5)
4.2 Similarity-based diverse tokens
Considering that attention in visual encoders often focuses more on foreground objects [6, 13], to prevent performance degradation on background-related questions after pruning, we also retain a set of diverse tokens as a complement. Among the remaining tokens, we first calculate the cosine similarity between each token as a metric of its diversity. Then, we gradually remove tokens with the highest similarity to others until the number of tokens left meets the requirement $R\cdot(1-r)$, implemented in Algorithm 1.
These diverse tokens ${\bm{Z}}_{v}^{div}$ better cover more areas of the input image and complement the important tokens, thus preserving as much information in the input image as possible.
Algorithm 1 Similarity-based Duplication Removal
Input: normalized: normalized visual tokens (N, C), n: diverse token number, r: removal number each iteration Output: remaining_idx: index of diverse tokens
while r > 0:
remaining = normalized[remaining_idx]
a, b = remaining[::2], remaining[1::2]
score = a @ b.transpose(-1, -2)
score = score.max(dim=-1).values
diverse_idx = score.argsort(dim=-1, descending=True)[:, r:]
remaining_idx = cat(remaining_idx[::2][diverse_idx], remaining_idx[1::2], dim=-1)
r = min(r, remaining_idx.shape[0] - n)
4.3 Inference with visual pruning
With the attention-based important visual tokens ${\bm{Z}}_{v}^{imp}$ and the similarity-based diverse visual tokens ${\bm{Z}}_{v}^{div}$, we concatenate these two parts and sort them according to their original positions in the image patch sequence. The merged visual tokens, after passing through a multimodal projector $g$, are then combined with the language instruction ${\bm{H}}_{q}$ and fed into the language model $f_{\phi}$ to generate the response ${\bm{X}}_{a}$:
${\bm{H}}_{v}=g({\bm{Z}}_{v}^{imp}\oplus{\bm{Z}}_{v}^{div}),\quad{\bm{X}}_{a}=f_{\phi}({\bm{H}}_{v}\oplus{\bm{H}}_{q})$ (6)
Here, $\oplus$ denotes the concatenation operation. Thanks to the complementary design between the important tokens and diverse tokens, our VisPruner is able to maintain considerable performance even at high reduction ratios, significantly enhancing the inference efficiency of VLMs.
@@ 5 Experiments
In this section, we validate our method across various VLM architectures on comprehensive multi-modal tasks, including high-resolution image and video understanding. We compare our approach with multiple existing methods and conduct an ablation study and an efficiency analysis.
@@ 5.1 Experimental setup
Datasets. We conduct extensive experiments on 10 image-based multi-modal benchmarks, including common visual question answering tasks such as VQAv2 [16], GQA [22], VizWiz [17], ScienceQA-IMG [38], and TextVQA [46], as well as other multi-modal benchmarks such as POPE [30], MME [15], MMBench [37], MMBench-CN [37] and MM-Vet [61]. Additionally, we also experiment on 4 widely used video question answering benchmarks, including TGIF-QA [23], MSVD-QA [56], MSRVTT-QA [56], and ActivityNet-QA [62]. All experiments on these benchmarks follow the default settings and evaluation metrics. Details of each task are provided in the supplementary materials.
Model architectures. We apply VisPruner to various VLM architectures, including the LLaVA series such as LLaVA-1.5 [34], LLaVA-NeXT [35] for high-resolution inputs, and Video-LLaVA [32] for video understanding, as well as other widely used models like Qwen-VL [3], InternVL [10], and CogVLM [54]. For all models, we follow the same inference settings as the original papers.
Comparison methods. We choose ToMe [4], FastV [9], SparseVLM [65], LLaVA-PruMerge [45] and the latest VisionZip [59] as comparison methods, covering a range of pruning approaches across different stages within VLMs.
@@ 5.2 Main results
Method $\text{VQA}^{\text{V2}}$ GQA VizWiz $\text{SQA}^{\text{IMG}}$ $\text{VQA}^{\text{Text}}$ POPE MME MMB $\text{MMB}^{\text{CN}}$ MMVet Acc. Rel.
Upper Bound, All 576 Tokens ($\mathbf{100\%}$)
LLaVA-1.5-7B 78.5 62.0 50.0 66.8 58.2 85.9 1510.7 64.3 58.3 31.1 63.1 100.0%
Retain 128 Tokens ($\downarrow\mathbf{77.8\%}$)
ToMe [4](ICLR23) 63.0 52.4 50.5 59.6 49.1 62.8 1088.4 53.3 48.8 27.2 52.1 83.9%
FastV [9](ECCV24) 61.8 49.6 51.3 60.2 50.6 59.6 1208.9 56.1 51.4 28.1 52.9 85.4%
SparseVLM [65](ICML25) 73.8 56.0 51.4 67.1 54.9 80.5 1376.2 60.0 51.1 30.0 59.4 94.4%
PruMerge+ [45](2024.05) 74.7 57.8 52.4 67.6 54.3 81.5 1420.5 61.3 54.7 28.7 60.4 95.8%
VisionZip [59](CVPR25) 75.6 57.6 52.0 68.9 56.8 83.2 1432.4 62.0 56.7 32.6 61.7 98.4%
VisPruner(Ours) 75.8 58.2 52.7 69.1 57.0 84.6 1461.4 62.7 57.3 33.7 62.4 99.6%
Retain 64 Tokens ($\downarrow\mathbf{88.9\%}$)
ToMe [4](ICLR23) 57.1 48.6 50.2 50.0 45.3 52.5 922.3 43.7 38.9 24.1 45.7 73.9%
FastV [9](ECCV24) 55.0 46.1 50.8 51.1 47.8 48.0 1019.6 48.0 42.7 25.8 46.6 75.9%
SparseVLM [65](ICML25) 68.2 52.7 50.1 62.2 51.8 75.1 1221.1 56.2 46.1 23.3 54.7 86.4%
PruMerge+ [45](2024.05) 67.4 54.9 52.9 68.6 53.0 77.4 1198.2 59.3 51.0 25.9 57.0 90.6%
VisionZip [59](CVPR25) 72.4 55.1 52.9 69.0 55.5 77.0 1365.6 60.1 55.4 31.7 59.7 95.6%
VisPruner(Ours) 72.7 55.4 53.3 69.1 55.8 80.4 1369.9 61.3 55.1 32.3 60.4 96.6%
Retain 32 Tokens ($\downarrow\mathbf{94.4\%}$)
ToMe [4](ICLR23) 46.8 43.6 51.3 41.4 38.3 39.0 828.4 31.6 28.1 17.3 37.9 61.4%
FastV [9](ECCV24) 43.4 41.5 51.7 42.6 42.5 32.5 884.6 37.8 33.2 20.7 39.0 64.1%
SparseVLM [65](ICML25) 58.6 48.3 51.9 57.3 46.1 67.9 1046.7 51.4 40.6 18.6 49.3 77.9%
PruMerge+ [45](2024.05) 54.9 51.1 52.8 68.5 50.6 70.9 940.8 56.8 47.0 21.4 52.1 83.0%
VisionZip [59](CVPR25) 67.1 51.8 52.9 68.8 53.1 68.7 1247.4 57.7 50.3 25.5 55.8 89.0%
VisPruner(Ours) 67.7 52.2 53.0 69.2 53.9 72.7 1271.0 58.4 52.7 28.8 57.2 91.5%
Table 1: Performance comparison of different pruning methods on LLaVA-1.5-7B. Here, Acc. denotes the average accuracy across 10 benchmarks, and Rel. represents the average percentage of performance maintained at the corresponding reduction ratio.
We first apply VisPruner to the classic LLaVA-1.5 model and conduct a comprehensive comparison with existing methods. Tab. 1 shows the the performance of different pruning methods on the LLaVA-1.5-7B model when retaining only 128, 64, and 32 tokens. With 77.8% of the visual tokens pruned, VisPruner remarkably maintained almost all of the original performance. When the number of retained tokens is further reduced to 64, about one-tenth of the full sequence length, VisPruner only decreases the original performance by 3.4%, outperforming the best text-visual attention-based method, SparseVLM, by as much as 10.2%. It also beats the latest vision-based pruning method, VisionZip, which merges redundant visual tokens in addition to pruning. This result suggests that identifying diverse tokens is more crucial than simply merging similar ones. With only about 5% of the tokens retained, VisPruner is still able to maintain 91.5% of the performance, proving the effectiveness of its design that leverages both important tokens and diverse tokens as complements.
5.3 VisPruner with higher resolution
Method $\text{VQA}^{\text{V2}}$ GQA $\text{VQA}^{\text{Text}}$ POPE MME Acc. Rel.
Upper Bound, All 2880 Tokens ($\mathbf{100\%}$)
LLaVA-NeXT-7B 81.2 62.9 59.6 86.3 1513.8 73.1 100.0%
Retain 640 Tokens ($\downarrow\mathbf{77.8\%}$)
FastV [9](ECCV24) 78.9 60.4 58.4 83.1 1477.3 70.9 97.0%
SparseVLM [65](ICML25) 78.2 59.1 56.2 80.9 1456.3 69.4 94.9%
VisionZip [59](CVPR25) 79.2 60.1 58.5 82.2 1468.4 70.7 96.7%
VisPruner(Ours) 79.8 61.6 59.3 85.9 1480.7 72.1 98.6%
Retain 320 Tokens ($\downarrow\mathbf{88.9\%}$)
FastV [9](ECCV24) 71.9 55.9 55.7 71.7 1282.9 63.9 87.7%
SparseVLM [65](ICML25) 71.4 56.5 52.4 73.5 1342.7 64.2 87.9%
VisionZip [59](CVPR25) 74.2 58.1 55.3 75.0 1348.8 66.0 90.5%
VisPruner(Ours) 75.7 58.4 57.6 80.4 1370.1 68.1 93.3%
Retain 160 Tokens ($\downarrow\mathbf{94.4\%}$)
FastV [9](ECCV24) 61.8 49.8 51.9 51.7 1079.5 53.8 74.7%
SparseVLM [65](ICML25) 62.2 50.2 45.1 54.6 1167.1 54.1 74.5%
VisionZip [59](CVPR25) 67.3 54.3 54.7 59.4 1239.7 59.5 82.3%
VisPruner(Ours) 70.6 54.7 56.0 72.9 1226.0 63.1 86.7%
Table 2: Performance comparison of different pruning methods on LLaVA-NeXT-7B. Acc. denotes the average accuracy across benchmarks, and Rel. represents the average percentage of performance maintained at the corresponding reduction ratio.